1) There are four letters and four envelopes. If the letters are placed in the
envelopes at random, what is the probability that not more than one letter is placed in the correct envelope?
2) The equation of regression lines are
3x+12y=19
3y+9x=46
Find 1) Correlation co- efficient
2) Mean value of X and Y
3) Ratio of the co- efficient of variability of X to Y

Answer by stanbon(75887)

) There are four letters and four envelopes. If the letters are placed in the
envelopes at random, what is the probability that not more than one letter is placed in the correct envelope?
Probability that a letter is placed in the correct envelope = 1/4
Prob that "not more than one letter is correctly placed" means
0 or 1 are correctly placed.
Prob (0 are correctly placed) = (3/4)^4=0.3164
Prob (1 is correctly placed) = 4C1 (1/4)(3/4)^3=0.421875
Sine there are mutually exclusive events you add the probabilities
to get 0.73828125
Cheers,
Stan H.